### v2

# Active-Learning Inspired Ab Initio Theory-Experiment Loop Approach for Management of Material Defects: Application to Superconducting Qubits

Abstract
Surface oxides are associated with two-level systems (TLSs) that degrade the performance of niobium-based superconducting quantum computing devices. To address this, we introduce a predictive framework for selecting metal capping layers that inhibit niobium oxide formation. Using DFT-calculated oxygen interstitial and vacancy energies as thermodynamic descriptors, we train a logistic regression model on a limited set of experimental outcomes to successfully predict the likelihood of oxide formation beneath different capping materials. This approach identifies Zr, Hf, and Ta as effective diffusion barriers. Our analysis further reveals that the oxide formation energy per oxygen atom serves as an excellent standalone descriptor for predicting barrier performance. By combining this new descriptor with lattice mismatch as a secondary criterion to promote structurally coherent interfaces, we identify Zr, Ta, and Sc as especially promising candidates. This closed-loop strategy integrates first-principles theory, machine learning, and limited experimental data to enable rational design of next-generation materials.
Introduction. The rational design of materials has played a pivotal role in advancing various technologies by accelerating the development of materials with tailored properties. This approach has been extensively applied in fields ranging from fuel cell catalysis [1], energy storage materials [2], and photonic structures [3] to high performance polymers [4]. These advances have been enabled by computational tools, high-throughput screening, and increasingly, machine learning, which allow for the optimization of properties including bandgaps [5], thermal conductivity [6], and electrochemical activity [7]. Despite these advances, the vast majority of rational design efforts have focused on bulk properties of materials, with relatively little attention paid to defect energetics [8].
While much of rational materials design has focused on tuning bulk properties, defect energetics play a critical role in determining how atoms migrate, especially across thin films and interfaces. This is particularly important in the design of materials for quantum computing [9, 10], specifically, in superconducting quantum devices, where Josephson junctions (JJs) are formed by sandwiching thin insulating or oxidized layers between superconducting electrodes. In such systems, oxygen diffusion from adjacent capping layers or the ambient environment can lead to the formation of unwanted oxides at the interface, most notably niobium oxide [11, 12]. These oxides host two-level systems (TLSs) that couple to the electric field of the qubit, acting as parasitic resonators that degrade coherence times and limit device performance [13, 14]. Suppressing the formation of these oxides is therefore a key challenge in improving qubit performance. Various strategies in this regard have been explored, with some notable methods such as high-temperature heat treatment and etching seen to reduce TLS losses [11, 15, 16, 17]. One such method, namely encapsulation of the superconducting metal with a barrier surface, has also shown experimental promise [18]. However, such studies have been mostly experimental, with insufficient theoretical investigation exemplifying use of the Materials Genome Loop.
In this work, we introduce a predictive framework that leverages defect energies calculated from first principles to evaluate the formation of niobium oxide in superconducting qubit JJs. By quantifying key microscopic energies—specifically, vacancy formation energies in metal oxides and oxygen interstitial energies in metals, we construct a logistic regression model that captures the likelihood of oxide formation across a range of candidate metals. The result is a phase-space map that distinguishes between effective and ineffective oxygen barrier materials and enables us to correctly predict new successful barrier materials. More broadly, this work demonstrates how rational defect engineering can be used to design atomically coherent, oxidation-resistant interfaces for quantum devices.
Computational details. All ab initio density–functional theory (DFT) calculations were performed with the open-source plane-wave code JDFTx [19, 20]. Ultrasoft pseudopotentials from the GBRV library [21] were employed together with the Perdew–Burke–Ernzerhof functional revised for solids (PBEsol) [22]. Calculations used the generalized-gradient approximation (GGA) and periodic boundary conditions in all directions. Plane-wave kinetic-energy cut-offs were set to 20 hartree for Kohn–Sham wavefunctions and 100 hartree for the charge density.
For each metal-oxide composition we selected the structure with the lowest formation energy reported in the Materials Project database [23]. Supercells were then constructed so that every defect was separated from its periodic images by at least three metal–metal spacings in all Cartesian directions, suppressing spurious defect–defect interactions. The same supercells were used for pristine and defect calculations to ensure strict numerical consistency.
Brillouin-zone integrations employed Monkhorst–Pack $\mathbf{k}$-point meshes, offset by $(0.5,0.5,0.5)$, that were individually converged until the total-energy change was below $10^{-5}\,\text{hartree}$ ($\approx 0.27\,\text{meV}$). Electronic self-consistency cycles were deemed converged when the total-energy difference between successive iterations fell beneath the same $10^{-5}\,\text{hartree}$ threshold.
[…]
Logistic regression. Logistic regression is a supervised machine learning method well suited for modeling binary classification problems using continuous-valued descriptors [27]. In the context of this study, it is used to predict the probability that niobium oxide will form at an interface, based on defect energies of candidate barrier materials. This model is integrated into an active learning framework, where predictions are iteratively refined through a theory-experiment loop: candidate materials are recommended by the model, tested experimentally, and the resulting data are used to further train the model.
The model estimates the probability $P(x_{i}|\beta_{i})$ of oxide formation as a function of input features $x_{i}=\{x_{1},x_{2},...,x_{n}\}$ (e.g. defect formation energies) and fitting parameters $\beta_{i}=\{\beta_{0},\beta_{1},\beta_{2},...,\beta_{n}\}$. The prediction is made using a sigmoid function, which maps the weighted sum of descriptors to a probability between 0 and 1, and takes the following form:
$P(x_{i}|\beta_{i})=\frac{1}{1+\exp\left[-\left(\beta_{0}+\beta_{1}x_{1}+...+\beta_{n}x_{n}\right)\right]}$ (1)
where the coefficients $\beta_{i}$ are learned by minimizing the cross entropy between the predicted and experimental probabilities from the training data set. The cross entropy is defined as:
$H=-\sum_{i=1}^{N}\left[y_{i}\log P(x_{i}|\beta_{i})+(1-y_{i})\log(1-P(x_{i}|\beta_{i}))\right]$ (2)
where $y_{i}\in[0,1]$ denotes the experimental probability of oxide formation for sample $i$, and $N$ is the total number of training samples.
Because the model predicts the probability of niobium oxide formation, it must account for the inherent uncertainty in the experimental training data. To incorporate this, binary observations of whether niobium oxide formed or not were converted into soft labels that reflect the confidence level in each experimental outcome. Metals that were observed to permit oxide formation were assigned experimental probabilities around 0.9, based on the likelihood that the observed oxide originated from diffusion processes rather than fabrication defects. If sample quality was independently verified through direct inspection, an experimental probability of 1 was assigned. Conversely, materials that prevented oxide formation were assigned experimental probabilities of 0. This assignment reflects the fact that, regardless of uncertainties in fabrication conditions, the metal successfully blocked oxide formation. The logistic regression model was then trained by minimizing the cross-entropy between the predicted probabilities and these experimental probabilities, thereby incorporating experimental uncertainty into the learning process.
Defect descriptors. To predict niobium oxide formation beneath metal capping layers, relevant thermodynamic descriptors were identified based on the microscopic processes that govern oxygen diffusion toward the niobium substrate. These processes are:
  1. 1.
    Metal Oxide to Metal transfers: In this process, an oxygen atom leaves the metal oxide leaving behind a vacancy and joins the metal in an interstitial site. The change in energy associated with this is
    $\Delta E_{\mathrm{MO\rightarrow M}}=E(V_{\mathrm{O}}(\mathrm{M}_{n}\mathrm{O}_{m}))+E(\mathrm{O}_{I}(\mathrm{M})),$ (3)
    where $E(V_{\mathrm{O}}(\mathrm{M}_{n}\mathrm{O}_{m}))$ is the energy to form an oxygen vacancy in the oxide $\mathrm{M}_{n}\mathrm{O}_{m}$, and $E(\mathrm{O}_{I}(\mathrm{M}))$ is the energy to add an oxygen interstitial into a metal $\mathrm{M}$. The details of the reference energies used for determining interstitial and vacancy energies are described below.
  2. 2.
    Metal to Niobium transfers: In this process, an oxygen atom leaves an interstitial site in the metal and joins the niobium in an interstitial site. This process is important to begin the formation of the niobium oxide. The change in energy associated with this is
    $\Delta E_{\mathrm{M\rightarrow Nb}}=-E(\mathrm{O}_{I}(\mathrm{M}))+E(\mathrm{O}_{I}(\mathrm{Nb})).$ (4)
  3. 3.
    Metal to Niobium Oxide transfers: In this process, an oxygen atom leaves an interstitial site in the metal and fills in a vacancy in the niobium oxide. This process is important for completing the niobium oxide. The change in energy associated with this process is
    $\Delta E_{\mathrm{M\rightarrow NbOx}}=-E(\mathrm{O}_{I}(\mathrm{M}))-E(V_{\mathrm{O}}(\mathrm{Nb}_{n}\mathrm{O}_{m})).$ (5)
The quantities that vary among different capping metals in Equations 3, 4, and 5 are $E(V_{\mathrm{O}}(\mathrm{M}_{n}\mathrm{O}_{m}))$ and $E(\mathrm{O}_{I}(\mathrm{M}))$. These are the defect energies that will be used as input features in the logistic regression model.
To evaluate our input features ab initio, we compute the oxygen interstitial energy (with O2 as the reference for the chemical potential of O) as
$E(\mathrm{O}_{I}(\mathrm{M}))=E(\mathrm{M}_{N}\mathrm{O})-\left[E(\mathrm{M}_{N})+\frac{1}{2}E(\mathrm{O}_{2})\right],$ (6)
where $E(\mathrm{M}_{N}\mathrm{O})$ is the energy of a computational supercell with $N$ atoms of metal M and one O interstitial, $E(\mathrm{M}_{N})$ is the energy of a supercell with $N$ atoms of metal M, and $E(\mathrm{O}_{2})$ is the energy of an oxygen molecule, which ultimately cancels out in Equations 3, 4, and 5. Similarly, the oxygen vacancy energy is computed as
$E(V_{\mathrm{O}}(\mathrm{M}_{n}\mathrm{O}_{m}))=$ $\left[E(\mathrm{M}_{Nn}\mathrm{O}_{Nm-1})+\frac{1}{2}E(\mathrm{O}_{2})\right]$
$-E(\mathrm{M}_{Nn}\mathrm{O}_{Nm})$ (7)
where $E(\mathrm{M}_{Nn}\mathrm{O}_{Nm-1})$ is the energy of a supercell with $N$ unit cells of the oxide $\mathrm{M}_{n}\mathrm{O}_{m}$ minus one O atom, and $E(\mathrm{M}_{Nn}\mathrm{O}_{Nm})$ is the energy of a supercell with $N$ unit cells of the oxide $\mathrm{M}_{n}\mathrm{O}_{m}$.
Figure 1: Diffusion barrier performance versus diffusion barrier energy: barrier energy (position along horizontal axis), performance (green=good, red=poor) for Au, Pd, W, Ta, Mo, Al, Pt, and Zr from left to right. Data show no clear correlation between barrier energies and performance.
To evaluate whether kinetic limitations might influence oxide formation, we also calculated diffusion barriers for oxygen hopping between interstitial sites in each barrier metal. However, as shown in Figure 1, there is no correlation between the computed diffusion barriers and the experimentally observed ability of metals to suppress niobium oxide formation. This strongly suggests that thermodynamic factors, not kinetics, control the outcome.
To reinforce this conclusion, we estimated the characteristic diffusion time for an oxygen atom to traverse a 5 nm capping layer. To estimate the diffusion constants, we use the standard Arrhenius expression:
$D=\frac{1}{6}a^{2}\nu\,e^{-\tfrac{E_{d}}{k_{B}T}}$ (8)
where $a\sim 0.3$ nm is the typical jump distance, $\nu\sim 10^{13}$ Hz is a characteristic phonon frequency, and $E_{d}$ is the diffusion barrier energy. Then, using $L\sim 5$ nm as the barrier thickness, we can estimate the corresponding diffusion times:
$\tau=\frac{L^{2}}{2D}$ (9)
For a moderate barrier of $E_{d}=0.5$ eV at room temperature, we find $\tau\sim 40$ ms. Even for higher barriers of $E_{d}\sim 0.8$ eV, the timescale remains on the order of a few minutes. This analysis shows that oxygen diffusion is generally fast enough to occur within experimental time frames across nearly all of our metals, except for those with the very highest barriers. Yet, we also find examples of low-barrier materials that successfully prevent oxide formation, and, conversely, our material with a $\sim$1.5 eV barrier (a $\sim 10^{12}$ increase in timescale over 0.8 eV) still allows for oxide formation. This indicates that oxygen can access alternative pathways (e.g., grain boundaries) to traverse the barrier. All together, these observations suggest that the presence or absence of niobium oxide is not controlled primarily by kinetic limitations, but rather more so by the thermodynamic favorability of the oxygen transfer processes described above. We therefore use only the interstitial and vacancy formation energies as descriptors in our model.
Phase diagram.
Figure 2: Oxide vacancy energy versus metal interstitial energy for a set of metals. Darker green regions indicate increasingly favorable thermodynamic barriers to oxygen diffusion
Figure 2 shows a phase diagram of oxygen vacancy energy versus metal interstitial energy for various candidate metals. Our initial processes in Equations 3, 4, and 5 represent the sequential microscopic steps necessary to form niobium oxide. For a metal to be a good diffusion barrier, we would like each of these processes to be thermodynamically unfavorable ($\Delta E>0$). Each of these conditions corresponds to a linear inequality in our two-dimensional descriptor space, forming a half-plane boundary. In Figure 2, the shaded green regions indicate the zones where each individual process is suppressed. Where these regions overlap, the shading becomes darker, indicating that multiple (or all) oxygen transfer pathways are energetically blocked. Metals located in these darker regions are therefore expected to serve as more effective diffusion barriers.
Most metals explored lie above the diagonal boundary corresponding to Equation 3, which prevents oxygen from leaching out of the surface oxide and entering the barrier metal. The vertical boundary near $E(\mathrm{O}_{I}(\mathrm{M}))\sim-3.5$ eV (from Equation 4) excludes metals that would readily transfer interstitial oxygen into metallic niobium as an interstitial. A stricter boundary near $E(\mathrm{O}_{I}(\mathrm{M}))\sim-6$ eV (from Equation 5) further excludes metals that would facilitate oxygen interstitial transfer into pre-existing niobium oxide vacancies. Both of these vertical cuts suggest that moving leftward in the plot corresponds to increasing thermodynamic resistance to oxide formation. Together, these three conditions indicate that materials in the upper-left quadrant of the phase space are most favorable, while those in the lower-right are least favorable.
At this level of analysis, of the potential barrier materials considered so far, one in particular, zirconium, stands out as satisfying all three criteria and therefore is a particularly promising candidate as a barrier material. Moreover, even if the zirconium is fully oxidized, the combination of Equations 3 and 4 ensures that it will maintain a sharp interface with niobium and thus zirconium oxide is also a particularly promising candidate as an insulating layer for Josephson junctions.
Figure 3: Oxygen vacancy energy vs. metal interstitial energy across logistic regression iterations: first iteration (a), final iteration (b). Experimentally tested metals are outlined in green (success) or red (failure). Point colors reflect predicted oxide formation probabilities. The solid blue line shows the model’s decision boundary; the dashed line indicates constant oxide formation energy.
Logistic theory-experiment loop. We note that the thermodynamic criteria discussed in the previous section, while physically motivated, may be too strict. Each of the three processes was treated independently, and the overall conclusion required all of them to be simultaneously unfavorable. However, in practice, preventing even a single step in the oxygen transfer pathway may be sufficient to block niobium oxide formation. Furthermore, the prior analysis does not account for competing effects (such as kinetics) that may impact the outcome over finite time scales. To move beyond a binary thermodynamic filtering, we adopt a data-driven approach based on logistic regression, which provides a probabilistic model trained directly on experimental results.
Our initial experimental data confirmed the broad trends predicted by the descriptor phase space, with successful barrier metals appearing toward the upper left of the phase space and unsuccessful metals appearing to the lower right. We then used this data to train the first iteration of our logistic regression model as described above in Logistic regression. Figure 3a shows the results, where experimental outcomes for Al, Ta, TiW3, Pd, Pt, and Au are marked by green or red circles outlining the data points, with green and red indicating whether niobium oxide formation was observed to be suppressed or not. These data were then used to train an initial logistic regression model, which outputs a continuous probability of oxide formation as a function of defect energies. The predicted probabilities are encoded in the figure as the color of the inner solid circles, ranging from green ($P=0$) to red ($P=1$).
Metals were then prioritized for further testing based on their predicted probability of success. After each new experiment, the result was fed back into the model, which was retrained to improve its predictive accuracy. This created a closed theory-experiment loop that progressed over several iterations. By the most recent iteration, shown in Figure 3b, the model had successfully predicted the behavior of five new potential barrier materials: Zr, Hf, Ti, Mo, and Ir, with one unsuccessful case: W. The final model shows strong consistency with experiment and captures smooth, interpretable probability gradients across the descriptor space.
Controlling process. The clean separation of data in Figure 3 by a single line suggests the existence of a single underlying controlling parameter that determines whether niobium oxide forms. Since our logistic regression model operates in a two-dimensional space of thermodynamic descriptors, the decision boundary it learns corresponds to a linear combination of the interstitial and vacancy formation energies. The line corresponding to a 50% probability of oxide formation provides the relation
$0.56E(\mathrm{O}_{I}(\mathrm{M}))-0.44E(V_{\mathrm{O}}(\mathrm{M}_{n}\mathrm{O}_{m}))=-1.36~\text{eV}.$ (10)
Physically, this combination can be interpreted as a process that introduces roughly 0.56 of an oxygen interstitial into the metal and fills in 0.44 of a vacancy in the metal oxide. One way to interpret this is as formation of a fraction of a metal oxide unit: oxygen atoms are first inserted into the metal as interstitials, and any remaining vacancies are subsequently eliminated. Quantitatively, this linear combination is reminiscent of an oxide formation process—specifically, the formation of one formula unit of the metal oxide from metal and oxygen atoms normalized by the number of oxygen atoms.
Figure 4: Per-oxygen oxide formation energy plotted against oxygen vacancy and metal interstitial energies for all calculated metals. Point colors indicate predicted oxide formation probabilities from the logistic regression model. A best-fit plane is shown, viewed at an angle that nearly reduces it to a line.
To explore this idea more directly, we examined the formation energy of metal oxides per oxygen atom, $E(\mathrm{M}_{n}\mathrm{O}_{m})/m$ and compared it to our original descriptors. Figure 4 plots the per-oxygen formation energy against the metal interstitial and oxide vacancy energies for each material that appears in Figure 3b. All materials are found to lie close to a single plane in this three-dimensional space, suggesting a tight linear relationship between these quantities. Accordingly, we fit these data with a least-squares plane, resulting in:
$E(\mathrm{M}_{n}\mathrm{O}_{m})/m=(0.19\pm 0.05)E(\mathrm{O}_{I}(\mathrm{M}))\\ -(0.77\pm 0.04)E(V_{\mathrm{O}}(\mathrm{M}_{n}\mathrm{O}_{m}))$ (11)
These coefficients correspond to the insertion of approximately 0.19 interstitials and removal of approximately 0.77 vacancies per oxygen atom (for a total insertion of 0.96 atoms), closely aligning with the stoichiometry of a typical metal oxide formula unit normalized to one oxygen atom. This strongly supports the interpretation that our two original descriptors are simply two linear components of the per-oxygen oxide formation energy.
To test whether the per oxygen oxide formation energy is sufficient to explain our logistic regression predictions, we next project a line of constant oxide formation energy (determined by logistical regression holding the slope of the line fixed) back onto the original descriptor phase space, as seen in Figure 3b. The resulting line separates the data just as well as the decision boundary learned by our original logistic regression model, indicating that per-oxygen oxide formation energy alone is sufficient to reproduce the model’s classification performance.
Figure 5: Per-oxygen oxide formation energy of various experimentally tested oxides, colored according to their success (green) or failure (red) to prevent oxidation of the niobium layer
Finally, to demonstrate the performance of the per oxygen oxide formation energy as a single descriptor, we plot this quantity for each experimentally tested metal along a single axis (Figure 5), coloring each point according to the observed oxygen barrier performance. A clear separation emerges between effective and ineffective barrier materials. This confirms that oxide formation energy per oxygen atom captures the dominant physics: if forming the oxide is thermodynamically favorable, the metal will tend to prevent oxidation of the buried niobium interface.
Coherent interfaces.
Figure 6: Oxide formation energy versus lattice mismatch percentage against niobium: Metals have been colored according to their logistic predictions, along with outer circles indicating experimental results denoted by their color
Having established that oxide formation energy is the primary thermodynamic descriptor governing niobium oxide suppression, we next consider the structural compatibility between candidate barrier metals and niobium. Even materials that thermodynamically inhibit oxide formation may introduce defects if they form incoherent interfaces due to poor lattice matching. Such mismatches can generate dislocations and grain boundaries that act as pathways for oxygen diffusion. To assess this, we computed the lattice mismatch percentage between each candidate metal and BCC niobium, and plotted it against oxide formation energy, as shown in Figure 6.
For face-centered cubic (FCC) metals, the mismatch was calculated using the FCC lattice constant divided by $\sqrt{2}$, corresponding to the in-plane spacing along the interface surface. For hexagonal close-packed (HCP) metals, we used the lattice constant of their lowest-energy FCC polymorph, consistent with the expectation that epitaxial thin films will favor FCC growth over BCC Nb.
We again observe a clear vertical separation between poor and effective oxygen barriers based on oxide formation energy, while the horizontal axis (lattice mismatch) distinguishes good materials from potentially optimal ones. Notably, Hf, Zr, Sc, and Ta exhibit both low oxide formation tendencies and reasonably small mismatch percentages, making them ideal candidates for coherent, oxidation-resistant interfaces. In contrast, metals such as Al and Ti show much higher mismatch values, suggesting they may still be effective as oxygen barriers but could exhibit greater variability or interface instability. From this perspective, we would rank Zr, and potentially Ta, actually as superior to Hf, and Sc (apart from its expense) as a particularly promising candidate that we have not yet explored experimentally. Together, these two material descriptors provide a simple yet powerful framework for identifying barrier materials that balance chemical stability with structural coherence.
Conclusion. In this work, we presented a data-driven framework that integrates DFT defect energetics with logistic regression to design capping layers suppressing niobium oxides in superconducting devices. Two physically meaningful descriptors—oxygen interstitial energies in metals and oxygen-vacancy energies in their oxides—define a learned decision boundary that separates effective from ineffective barriers, prospectively identifying Zr, Hf, and Ta, which we then verified experimentally. Adding lattice mismatch as a secondary screening criterion prioritizes coherent interfaces. Within this framework, Zr and Ta combine low oxide propensity with small mismatch, Hf proves effective albeit more mismatched, and Sc emerges as a new potential candidate.
This work highlights a generalizable methodology for integrating defect-level energetics into the rational design of functional materials. More broadly, the strategy is applicable whenever readily computable, physically controlling descriptors can be identified. By unifying such descriptors with limited experimental studies through carefully selected machine learning techniques, as demonstrated here, we establish a robust closed-loop materials-by-design framework that accelerates the discovery of next-generation materials.
Acknowledgments. This work was supported by the Air Force Office of Scientific Research under award number FA9550-23-1-0706. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the United States Air Force. This work made use of the Cornell Center for Materials Research shared instrumentation facility for the XPS analysis system.
Data availability. The data that support the findings of this article are openly available [28].